Help solving these problems
a for-profit hospital, has $1 million END-OF-CHAPTER PROBLEM in taxabe income for 2016, and its tax rate is 30 percent. a. Given this information, what is the firms net income2 (Net income is what remains after taxes have been paid.) b. Suppose the hospital pays out $300,000o in divide areholder, Carl n dividends is 15 percent, what is his receives St0,00o. If Carls tax rate on dividends is the stock of another corporation does not have to pavts nt of dividends received. In general, only 30 percent of thed tax dividend? taxes on the A firm that owns 2.2 this tax law ire amount o en received by one corporation from another are taxable. The purpose o feature is to mitigate the effect of triple taxation, which occurs when first taxed at the first firm, its dividends paid to the second firm are taxed5e and the dividends paid to shareholders by the second firm are tax gain, t again Assume that a firm with a 21 percent tax rate receives S after-tax amount of the dividend? ing in HCA (taxable) bonds that carry a 12 percent interest rate. 100,000 in dividend s f t is th another corporation. What taxes must be paid on this dividend, and wha 2.3 Theresa Davis is in the 4o percent personal tax bracket. She is considering invest a. What is her after-tax yield (interest rate) on the bonds? Suppose Twin Cities Memorial Hospital has issued tax-exempt bonds that hae an interest rate of 6 percent. With all else the same, should Theresa buy the HCA or the Twin Cities bonds? With all else the same, what interest rate on the tax-exempt Twin Cities bonds c. would make these bonds and the HCA bonds equally advantageous? 2.4 Aditi Patel currently holds tax-exempt bonds of Good Samaritan Healthcare that pay 7 percent interest. She is in the 40 percent tax bracket. Her broker wants hert buy some Beverly Enterprises taxable bonds that will be issued next week else the same, what rate must be set on the Beverly bonds to make Aditi in making a switch? With all iti interested eorge and Margaret Wealthy are in the 48 percent tax bracket, considerings federal and state personal taxes. Norman Briggs, the CEO of com Hospital, has been aggressively pursuing a contribution from the the hospitals soon-to-be-built Cancer Care Center. Without the con munit of $500,000 to contribution, the althystaxable income for 2016 would be $2 million. What impadt contribution have on the Wealthys 2016 tax bill?
Answer #1

(2.1) (a) Taxable Income = $ 1000000 and Tax Rate = 30 %

Net Income = Taxable Income - Tax Expense = 1000000 - 0.3 x 1000000 = $ 700000

(b) Dividends Received = $ 10000 and Tax Rate on Dividends = 15 %

After-Tax Dividend = Dividends Received - Tax on Dividends = 10000 - 0.15 x 10000 = $ 8500

(2.2) Dividends Received = $ 100000, Of the dividends received only 30 % is taxable at a rate of 21 %

Tax Expense = 0.21 x 0.3 x 100000 = $ 6300

After-Tax Dividend = 100000 - 6300 = $ 93700

(2.3) (a) Interest Rate = 12 % and Tax Rate = 40 %

After-Tax Yield = 12 x (1-0.4) = 7.2 %

(b) The tax exempt bonds provide an interest of 6 % whereas the after-tax yield of the HCA bonds are higher at 7.2 %

Therefore, the investor will go with the higher interest bonds which is the HCA

(c) The tax exempt bonds should have an interest rate equal to the HCA bond's after-tax yield so as to bring both investments at par and have the same advantage.

(2.4) Curent Bond Interest Rate = 7 % and Tax Rate = 40 %

In order for the new bonds to be at par with the current bonds, the formers after-tax yield needs to be equal to the current bond's interest rate

Therefore, R x (1-0.4) = 7 % where R is the new bond's pre-tax yield

R = 7 / (1-0.4) = 11.67 %

NOTE: Please raise separate queries for solutions to the remaining unrelated questions, as one query is limited to the solution of only one question with a maximum of four sub-parts.